! Required information Gable Company uses three activity cost pools. Each pool has a cost driver. Information for Gable CompaRequired 1 Required 2 Classify each activity as facility, product, batch, o Machining Design SetupSuppose that Gable Company manufactures three products, A, B, and C. Information about these products follows: Number of machSuppose that Gable Company manufactures three products, A, B, and C. Information about these products follows: Number of mach
Answer #1

Gable Company

1.

Activity Activity Rate
Machining $              4.50 per Machine Hour
Design Costs $              9.80 per Design Hour
Setup $          143.80 per Batch

2.
Machining - Unit level
Design Costs - Product level
Setup - Batch level

Overhead assigned
Product A = 36000 x $4.50 + 3800 x $9.80 + 50 x $143.80 = $206430
Product B = 46000 x $4.50 + 2100 x $9.80 + 175 x $143.80 = $252745
Product C = 10600 x $4.50 + 3600 x $9.80 + 285 x $143.80 = $123963

Activity Proportion
Overhead assigned
Product A = $206430
Product B = $252745
Product C = $123963

If you have any query, kindly comment with your query and please mark thumbs up.